# Towards Resilient and Efficient LLMs:

### v2

@@ Abstract
With the increasing demand for practical applications of Large Language Models (LLMs), many attention-efficient models have been developed to balance performance and computational cost. However, the adversarial robustness of these models remains under-explored. In this work, we design a framework to investigate the trade-off between efficiency, performance, and adversarial robustness of LLMs by comparing three prominent models with varying levels of complexity and efficiency – Transformer++, Gated Linear Attention (GLA) Transformer, and MatMul-Free LM – utilizing the GLUE and AdvGLUE datasets. The AdvGLUE dataset extends the GLUE dataset with adversarial samples designed to challenge model robustness. Our results show that while the GLA Transformer and MatMul-Free LM achieve slightly lower accuracy on GLUE tasks, they demonstrate higher efficiency and either superior or comparative robustness on AdvGLUE tasks compared to Transformer++ across different attack levels. These findings highlight the potential of simplified architectures to achieve a compelling balance between efficiency, performance, and adversarial robustness, offering valuable insights for applications where resource constraints and resilience to adversarial attacks are critical.
@@ I Introduction
In this paper, we aim to fill the research gap by presenting a framework to assess trade-offs between computational efficiency, performance, and adversarial robustness of LLMs and applying our framework to three prominent models with different complexity – Transformer++ [15], GLA Transformer [11], and MatMul-Free LM [12]. Transformer++ is an advanced architecture known for its high performance, while GLA Transformer and MatMul-Free LM are recent attention-efficient models with innovative attention mechanisms and proven computational efficiency. In our framework, we perform task-specific fine-tuning and evaluation of the three models on four NLP classification tasks from the GLUE dataset [16]. Then we evaluate the fine-tuned models on the corresponding tasks from the AdvGLUE dataset [17], which includes word-level, sentence-level, and human-level adversarial attacks to the GLUE dataset. Our experiments show that GLA Transfomer and MatMul Free LM achieve higher efficiency and comparative performances compared to Transformer++ across GLUE tasks. In addition, GLA Transfomer demonstrates superior robustness across all attack levels, while MatMul-Free LM is more robust to word-level attacks and equally robust to sentence-level and human-level attacks as Transformer++. These findings offer valuable insights into the applicability and resilience of these models, thereby informing future development and deployment strategies in different environments.
@@
    We introduce a framework and present empirical studies to assess trade-offs between computational efficiency, performance, and adversarial robustness of LLMs with varying complexity.
@@ II-A Vulnerability of LLM
Large Language Models (LLMs) have demonstrated remarkable performances across a variety of tasks, driving revolutionary changes and developments in various domains where reliability and robustness are critical. For example, in the medical fields, LLMs have been employed to streamline the process of determining patient eligibility for clinical trials [18], facilitate drug information management through conversational interfaces [19], and predict chronic disease [20]. In software security applications, LLMs have been integrated with traditional fuzzing techniques to improve software vulnerability detection [21], trained to enhance the code review process and detect security issues in software [22]. Failures in such domains can lead to severe consequences, raising significant concerns about the reliability of LLMs. Despite their advanced capabilities, LLMs and related applications have been shown to possess vulnerabilities. For instance, Wallace et al. unveiled that subtle yet sophisticated modifications from human-generated adversarial examples can significantly mislead models to generate incorrect responses [23]. Retrieval-based Augmented Generative Models (RAG), which enhance LLMs by integrating external knowledge, can be manipulable if malicious content is ingested into their knowledge bases [24]. To mitigate the vulnerabilities of LLMs, ongoing studies have been investigating various solutions, including but not limited to incorporating adversarial training techniques [25], adopting meta-learning techniques [26] to adapt defenses against adversarial attacks, systematically rectifying LLMs via dead-end analysis[27], and integrating uncertainty-aware models to handle ambiguous inputs [28]. However, existing approaches still fail to provide comprehensive protection against evolving forms of vulnerabilities, highlighting the need for ongoing research to address a wide range of threats and ensure the robustness of LLMs in diverse and dynamic environments.
@@ III-A E-P-R Trade-off Evaluation Framework
In this work, we present an E-P-R Trade-off Evaluation Framework, as illustrated in Figure Figure 1, to assess the trade-offs between efficiency, performance, and adversarial robustness of Large Language Models (LLMs). Our framework begins by fine-tuning pre-trained LLMs on a variety of language tasks. Subsequently, various adversarial attacks, such as word-level, sentence-level, and human-level perturbations, are applied to or added to original samples from each language task to generate adversarial examples. The fine-tuned models are evaluated for their performance and adversarial robustness utilizing the original and adversarial samples. Additionally, the fine-tuning or inference efficiency of each LLM is assessed. This holistic approach provides a detailed understanding of LLM capabilities and the trade-offs among efficiency, performance, and adversarial robustness.
@@ III-B Large Language Models
In this work, we focus on three language models: Transformer++ [15], GLA Transformer [11], and MatMul Free LM [12]. Each of the three models employs different attention mechanisms, resulting in varying computational complexities and efficiencies. These models have demonstrated comparative performances on various language tasks with different training efficiency.
Transfomer++ Transformer++ is an advanced architecture that builds upon the original transformer architecture [1]. Transformer++ follows the encoder-decoder architectures of Transformer while introducing a new form of multi-head attention that learns in-context dependencies through convolution. Specifically, Transformer++ employs convolution-based attention by using an adaptive sequence module and adopts an adaptive query module and a dynamic convolution module to capture local word context and entire sequence context, respectively. Given the total number of multi-head (H) and weight matrices $W_{i}^{Q},W_{i}^{K}$, and $W_{i}^{V}$, Equation (1) shows how the model combines convolution-based attention with traditional self-attention that captures global dependencies to form a more comprehensive attention mechanism [13]. Transformer++ has been demonstrated to be a strong transformer baseline model for various language tasks in previous studies.
@@ III-C Dataset
We utilize the General Language Understanding Evaluation (GLUE) [16] and Adversarial GLUE (AdvGLUE) [17] benchmark datasets to evaluate the performance and adversarial robustness of LLMs.
GLUE is a widely recognized framework that evaluates language models across a diverse set of natural language processing (NLP) tasks, including sentiment analysis, paraphrase detection, and sentence similarity.
AdvGLUE, on the other hand, extends the GLUE benchmark by introducing 14 distinct adversarial attack methods, providing a rigorous test for model robustness against adversarial inputs. The AdvGLUE benchmark is essential for understanding how models perform not just under standard conditions but also in the presence of adversarial challenges. AdvGLUE employs the same evaluation metrics as GLUE, enabling a direct and fair comparison between performance on benign and adversarial examples.
In our study, we evaluate the performance and adversarial robustness of LLMs using four key tasks: Standard Sentiment Analysis (SST-2), Multi-Genre Natural Language Inference (MNLI), Quora Question Pairs (QQP), and Question Natural Language Inference (QNLI). These tasks were chosen for their relevance and ability to test various aspects of language understanding and inference.
@@ V Results and Discussion
MODEL SST-2 QQP QNLI MNLI AVG SST-2 QQP QNLI MNLI AVG
@@
Human-level Word-level Sentence-level AVG
@@ VI Conclusion
There are several limitations in our work. While AdvGLUE encompasses a variety of language tasks and diverse adversarial attack methods, these attacks are designed to target known vulnerabilities only. With more vulnerabilities uncovered, future research should broaden evaluations to explore subtle differences in robustness. Meanwhile, due to resource constraints, only selected LLMs of a specific size was evaluated. Future work should apply the framework to models of various complexities and sizes to obtain a more comprehensive analysis. Additionally, our work primarily evaluates model performance and adversarial robustness in a controlled experimental setting. Real-world applications may present other challenges not captured in our study, such as handling noisy or incomplete data, integration with other systems, and latency concerns. Future research should explore these practical considerations in more diverse operational contexts.

### v3

@@ Abstract
With the increasing demand for practical applications of Large Language Models (LLMs), many attention-efficient models have been developed to balance performance and computational cost. However, the adversarial robustness of these models remains under-explored. In this work, we design a framework to investigate the trade-off between efficiency, performance, and adversarial robustness of LLMs and conduct extensive experiments on three prominent models with varying levels of complexity and efficiency – Transformer++, Gated Linear Attention (GLA) Transformer, and MatMul-Free LM – utilizing the GLUE and AdvGLUE datasets. The AdvGLUE dataset extends the GLUE dataset with adversarial samples designed to challenge model robustness. Our results show that while the GLA Transformer and MatMul-Free LM achieve slightly lower accuracy on GLUE tasks, they demonstrate higher efficiency and either superior or comparative robustness on AdvGLUE tasks compared to Transformer++ across different attack levels. These findings highlight the potential of simplified architectures to achieve a compelling balance between efficiency, performance, and adversarial robustness, offering valuable insights for applications where resource constraints and resilience to adversarial attacks are critical.
@@ I Introduction
To the best of our knowledge, no comprehensive evaluation framework exists to assess the adversarial robustness of attention-efficient models. Previous studies have predominantly focused on performance and robustness of traditional transformer architectures, with little consideration of lightweight models designed for efficiency [15, 16, 17]. In this paper, we aim to fill the research gap by presenting a framework to assess trade-offs between computational efficiency, performance, and adversarial robustness of LLMs and applying our framework to three prominent models with different complexity – Transformer++ [18], GLA Transformer [11], and MatMul-Free LM [12]. Transformer++ is an advanced architecture known for its high performance, while GLA Transformer and MatMul-Free LM are recent attention-efficient models with innovative attention mechanisms and proven computational efficiency. In our framework, we perform task-specific fine-tuning and evaluation of the three models on four NLP classification tasks from the GLUE dataset [19]. Then we evaluate the fine-tuned models on the corresponding tasks from the AdvGLUE dataset [20], which includes word-level, sentence-level, and human-level adversarial attacks to the GLUE dataset. Our experiments show that GLA Transfomer and MatMul Free LM achieve higher efficiency and comparative performances compared to Transformer++ across GLUE tasks. In addition, GLA Transfomer demonstrates superior robustness across all attack levels, while MatMul-Free LM is more robust to word-level attacks and equally robust to sentence-level and human-level attacks as Transformer++. These findings offer valuable insights into the applicability and resilience of these models, thereby informing future development and deployment strategies in different environments.
@@
    We introduce a novel framework and conduct the first empirical studies to assess trade-offs between computational efficiency, performance, and adversarial robustness of LLMs with varying complexity.
@@ II-A Vulnerability of LLM
Large Language Models (LLMs) have demonstrated remarkable performances across a variety of tasks, driving revolutionary changes and developments in various domains such as art [21], healthcare [22, 23, 24], software security [25, 26], etc., some of where reliability and robustness are critical. Failures in such domains can lead to severe consequences, raising significant concerns about the reliability of LLMs. Despite their advanced capabilities, LLMs and related applications have been shown to possess vulnerabilities. For instance, Wallace et al. unveiled that subtle yet sophisticated modifications from human-generated adversarial examples can significantly mislead models to generate incorrect responses [27]. Retrieval-based Augmented Generative Models (RAG), which enhance LLMs by integrating external knowledge, can be manipulable if malicious content is ingested into their knowledge bases [28]. To mitigate the vulnerabilities of LLMs, ongoing studies have been investigating various solutions, including but not limited to incorporating adversarial training techniques [29], adopting meta-learning techniques [30] to adapt defenses against adversarial attacks, systematically rectifying LLMs via dead-end analysis[31], and integrating uncertainty-aware models to handle ambiguous inputs [32]. However, existing approaches still fail to provide comprehensive protection against evolving forms of vulnerabilities, highlighting the need for ongoing research to address a wide range of threats and ensure the robustness of LLMs in diverse and dynamic environments.
@@ III-A E-P-R Trade-off Evaluation Framework
In this work, we propose the Efficiency-Performance-Robustness (E-P-R) Trade-off Evaluation Framework, as illustrated in Figure 1, to comprehensively assess Large Language Models (LLMs) based on three critical aspects: computational efficiency, task performance, and adversarial robustness. This framework provides a structured approach to understanding the trade-offs between these aspects, facilitating informed decision-making when deploying LLMs in real-world applications. The framework consists of three key components: task-specific fine-tuning, adversarial attack generation, and trade-off evaluation.
Task-specific Fine-tuning The first part is fine-tuning pre-trained LLMs on a variety of language tasks to adapt the models for specific applications. This step ensures that each model is optimized for the specific task while maintaining comparability across models. The fine-tuned models are prepared for evaluation under both standard and adversarial conditions.
Adversarial Attack To assess adversarial robustness, we introduce perturbations to the input data at various levels, including word-level, sentence-level, and human-level. These adversarial examples are generated by applying different types of perturbations to the original input samples from each task. The robustness of each LLM is then evaluated based on its performance on these adversarial examples, allowing for a detailed examination of how well the models handle malicious or unexpected input variations.
Trade-off Evaluation Once the models have been fine-tuned and tested under adversarial conditions, we conduct a comprehensive analysis of the trade-offs between efficiency, performance, and robustness. Efficiency is measured in terms of computational resources required for fine-tuning and inference, while performance and robustness are evaluated on both standard and adversarial data. The adversarial robustness of each model is assessed based on its ability to maintain performance under various levels of adversarial attacks. By examining the interplay between these three aspects, our framework enables a nuanced understanding of the strengths and limitations of each model, providing insights for optimizing model deployment in different environments.
@@ III-B Large Language Models
In this work, we focus on three language models: Transformer++ [18], GLA Transformer [11], and MatMul Free LM [12]. These models are carefully selected to represent distinct approaches to the core attention mechanism, which is central to transformer architectures and directly impacts model scalability and practicality across different application contexts. They have all demonstrated comparative performances on various language tasks with different training efficiency, but their adversarial robustness remains unexplored. Comparing these three models enables a holistic evaluation of the trade-offs between computational efficiency, performance, and adversarial robustness.
Transformer++ Transformer++ is an advanced architecture that builds upon the original transformer architecture [1]. Transformer++ follows the encoder-decoder architectures of Transformer while introducing a new form of multi-head attention that learns in-context dependencies through convolution. Specifically, Transformer++ employs convolution-based attention by using an adaptive sequence module and adopts an adaptive query module and a dynamic convolution module to capture local word context and entire sequence context, respectively. Given the total number of multi-head (H) and weight matrices $W_{i}^{Q},W_{i}^{K}$, and $W_{i}^{V}$, Equation (1) shows how the model combines convolution-based attention with traditional self-attention that captures global dependencies to form a more comprehensive attention mechanism [13]. Transformer++ has been demonstrated to be a strong transformer baseline model for various language tasks in previous studies.
@@ III-C Dataset
In this study, we utilize both the General Language Understanding Evaluation (GLUE) [19] and Adversarial GLUE (AdvGLUE) [20] benchmark datasets to comprehensively evaluate the performance and adversarial robustness of large language models (LLMs). These datasets are specifically chosen for their ability to test a wide range of natural language processing (NLP) skills and their relevance to real-world language understanding tasks.
GLUE is a widely recognized benchmark designed to evaluate the generalization capabilities of language models across a diverse set of NLP tasks. These tasks include sentiment analysis, paraphrase detection, and sentence similarity, among others, which collectively cover multiple facets of language understanding such as syntax, semantics, and inference. The primary reason for selecting GLUE is its well-established use as a standard benchmark for evaluating NLP models, which allows comparison with a broad set of results from the research community. GLUE’s inclusion ensures that our evaluation covers a comprehensive range of language tasks, ensuring that our results reflect the overall language processing capabilities of the models under study.
On the other hand, AdvGLUE extends the GLUE benchmark by introducing a set of 14 distinct adversarial attack methods that rigorously test model robustness against adversarial inputs. The rationale for incorporating AdvGLUE is to move beyond performance evaluation under standard conditions and assess how well the models perform when faced with adversarial challenges. Robustness to adversarial inputs is a critical consideration for real-world applications, especially in domains where language models might be deployed in environments prone to manipulation or adversarial attacks, such as automated customer service, security systems, or sensitive decision-making processes. AdvGLUE uses the same evaluation metrics as GLUE, allowing for a direct and fair comparison between performance on benign examples and adversarially perturbed inputs. This ensures a balanced view of both standard and adversarial model performance, a key focus of our research.
In our study, we focus on four key tasks from both GLUE and AdvGLUE: Standard Sentiment Analysis (SST-2), Multi-Genre Natural Language Inference (MNLI), Quora Question Pairs (QQP), and Question Natural Language Inference (QNLI). These tasks are chosen not only because of their relevance to various real-world applications but also because they cover different aspects of language understanding and inference. By evaluating models on both standard and adversarial versions of these tasks, we are able to thoroughly assess their performance and robustness.
@@ V Results and Discussion
SST-2 QQP QNLI MNLI AVG SST-2 QQP QNLI MNLI AVG
@@
Model Human-level Word-level Sentence-level AVG
@@ VI Conclusion
While our framework and experiment provide valuable insights, there are several limitations. First, the efficiency evaluation is limited to fine-tuning and inference times on specific hardware, which may not fully represent real-world environments with varied infrastructure and operational costs. Second, our adversarial robustness testing is confined to limited attack methods, leaving potential vulnerabilities from newer attack strategies unaddressed. Third, we test models of specific sizes due to resource limitations, which restricts the generalization of our results across different model scales. Additionally, the framework does not account for real-world challenges such as noisy or incomplete data, system latency, or integration complexities that could affect model performance and robustness. Future research should focus on broadening the adversarial attack spectrum, applying the framework to a wider range of model sizes and architectures, and incorporating real-world operational constraints to make the framework more applicable to diverse deployment environments.